1. Consider a perfectly competitive market with demand curve given by P, 200 D. The industry supply curve in this market is PsQs (a) Draw the demand-supply graph for this market. Calculate the quantit;y traded, equilibrium price for this market. Also calculate the Total Consumer Surplus (TCS) and Total Producer Surplus (TPS) for this market (b) Suppose that the government is considering a price ceiling, P1 - $20 Find the quantity traded, equilibrium price, TCS and TPS under the price ceiling to the corrsponding values in #lal. Give numerical solutions and show how you get your answer. (c) Suppose that the government is considering a price floor, and P2 $60 Find the quantity traded, equilibrium price, TCS and TPS under the price floor to the corrsponding values in #LalGive numerical solutions and show how you get your answer. d) What is the deadweight loss in?1b and 1clGive numerical solutions and show how you get your answer. (e) Suppose that the government is considering legislating a production quota of 35 units. How much should the government pay the producers to ensure that they produce exactly 35 units and no more? What is the TCS and TPS under this system? Give numerical solutions and show how you get your answer. (f) (**) 2 points Suppose that the government is considering imposing a sup- port price of $60/unit. How much will the government need to raise in taxes to ensure that it can buy the excess production at this support price? Give numerical solutions and show how you get your answer
Answer #1

Pb = 200 - 4Qd

Ps = Qs

At equilibrium Pb = Ps

200 - 4Q = Q

200 = 5Q

Q = 40

P = 40

Now the support price is 60

At the price of 60, the supply would be Qs = Ps = 60

At this price, the demand would be 60 = 200-4Qd = 35

Thus the government has to purchase the surplus of 60-35 = 25

Thus the government has to raise 25*60 = 1500.

please give a thumbs up if you found the answer helpful.
This is my solution
Thank you...